# Stealthy Attack on Large Language Model based Recommendation

### v1

@@ Abstract
Recently, the powerful large language models (LLMs) have been instrumental in propelling the progress of recommender systems (RS). However, while these systems have flourished, their susceptibility to security threats has been largely overlooked. In this work, we reveal that the introduction of LLMs into recommendation models presents new security vulnerabilities due to their emphasis on the textual content of items. We demonstrate that attackers can significantly boost an item’s exposure by merely altering its textual content during the testing phase, without requiring direct interference with the model’s training process. Additionally, the attack is notably stealthy, as it does not affect the overall recommendation performance and the modifications to the text are subtle, making it difficult for users and platforms to detect. Our comprehensive experiments across four mainstream LLM-based recommendation models demonstrate the superior efficacy and stealthiness of our approach. Our work unveils a significant security gap in LLM-based recommendation systems and paves the way for future research on protecting these systems.
@@
Jinghao Zhang1,2, Yuting Liu3, Qiang Liu1,2, Shu Wu1,2, Guibing Guo3, Liang Wang1,2, 1Center for Research on Intelligent Perception and Computing, State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, Chinese Academy of Sciences, 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Northeastern University, China {jinghao.zhang}@cripac.ia.ac.cn, {qiang.liu,shu.wu,wangliang}@nlpr.ia.ac.cn {yutingliu}@stumail.neu.edu.cn, {guogb}@swc.neu.edu.cn
@@ 1 Introduction
Contrary to traditional recommendation models, which rely heavily on abstract and less interpretable ID-based information, LLM-based recommendation models exploit the semantic understanding and strong transferability of LLMs. This approach places a heightened focus on the textual content of items, such as titles and descriptions Lin et al. (2023); Chen et al. (2023). For instance, many researchers Hou et al. (2022, 2023a); Yuan et al. (2023); Li et al. (2023a); Yang et al. (2023); Geng et al. (2022); Cui et al. (2022); Bao et al. (2023a); Zhang et al. (2023a); Li et al. (2023b); Zhang et al. (2023b) have explored modeling user preferences and item characteristics through a linguistic lens. This methodology promises a revolutionary shift in the conventional paradigm of recommendations by providing generalization capabilities to novel items and datasets.
@@
Despite these advancements, the security of RS remains a largely unaddressed issue. Malicious attacks on these systems can lead to undesirable outcomes, such as the unwarranted promotion of low-quality products in e-commerce platforms or the spread of misinformation in news dissemination contexts. Traditional shilling attack strategies on RS Wang et al. (2023a, 2024) involve the generation of fake users who are programmed to give high ratings to specific target items. By introducing such cheating data, it aims at influencing the training of the recommender models and subsequently increasing the exposure of the target items.
@@
We construct comprehensive experiments on four mainstream LLM-based recommendation models Geng et al. (2022); Bao et al. (2023a); Li et al. (2023a); Zhang et al. (2023b) as victim models to validate the outstanding efficacy and stealthiness of the textual attack paradigm compared with traditional shilling attacks Burke et al. (2005a); Kaur and Goel (2016a); Lin et al. (2020). We further delve into the effects of model fine-tuning and item popularity on the attack. Additionally, we investigate the transferability of the attack across various victim models and recommendation tasks to demonstrate its practical applicability and utility in real-world scenarios. Finally, we evaluate a simple re-writing defense strategy, which also can mitigate the issue to some extent.
@@ 3.1.1 Victim Models
We choose four mainstream LLM-based recommendation models as our victim models: Recformer Li et al. (2023a), P5 Geng et al. (2022), TALLRec Bao et al. (2023a) and CoLLM Zhang et al. (2023b). Please refer to Appendix A.1 for more details.
@@ 3.1.2 Compared Shilling Attacks
Shilling attacks aim to generate fake users that assign high ratings for a target item, while also rating other items to act like normal users for evading. We compare our text attack paradigm with white-box shilling attacks Random attack Kaur and Goel (2016a), Bandwagon attack Burke et al. (2005a), and gray-box Aush Lin et al. (2020) and Leg-UP Lin et al. (2024). Please refer to Appendix A.2 for more details.
@@ 3.1.3 Datasets
We conduct experiments on three categories of widely-used Li et al. (2023a); Geng et al. (2022); Bao et al. (2023a); Zhang et al. (2023b) Amazon review dataset introduced by McAuley et al. (2015): ‘Beauty’, ‘Toys and Games’, ’Sports and Outdoors’, which are named as Beauty, Toys and Sports in brief. We use the 5-core version of Amazon datasets where each user and item have 5 interactions at least. The statistics of these datasets are summarized in Appendix A.3.
@@ 3.4 The influence of item popularity
In this section, we examined the impact of the initial popularity of target items on attacks. Popularity bias in recommendation systems favors popular items over personalized ones, limiting diversity, fairness and potentially dissatisfying users’ preferences Zhang et al. (2021). We selected the top 150 and bottom 150 items in popularity from the dataset as target items and presented the attack results in Table 4. We can observe that items with high popularity experience greater promotion, thereby exacerbating popularity bias. High popularity target items can achieve greater exposure boosts with fewer queries and higher semantic consistency.
@@ 4.1 LLM-based Recommendation
The techniques used by LLMs in the recommendation domain involve translating recommendation tasks into natural language tasks and adapting LLMs to generate recommendation results directly. These generative LLM-based approaches can be further divided into two paradigms based on whether parameters are tuned: non-tuning and tuning paradigms. The non-tuning paradigm assumes LLMs already have the recommendation abilities and attempt to trigger the strong zero/few-shot abilities by introducing specific prompts Liu et al. (2023); Dai et al. (2023); Mysore et al. (2023); Wang et al. (2023b); Hou et al. (2023b). The tuning paradigm uses fine-tuning, prompt learning, or instruction tuning Kang et al. (2023); Bao et al. (2023b); Wang et al. (2022); Geng et al. (2022); Cui et al. (2022) to enhance LLM’s recommendation abilities by using LLMs as encoders to extract user and item representations and then fine-tuning their parameters on specific loss functions.
@@ 4.2 Shilling Attack
Shilling attacks aim to interfere with the recommendation strategy of a victim recommender system by injecting fake users into the training matrix Deldjoo et al. (2019); Toyer et al. (2023). This can be implemented through (1) heuristic attacks Burke et al. (2005b); Linden et al. (2003); Kaur and Goel (2016b), where fake profiles are created based on subjective inference and existing knowledge; (2) gradient attacks Fang et al. (2020); Li et al. (2016); Zhang et al. (2020); Fang et al. (2018); Huang et al. (2021), which optimize the objective function through a continuous space; and neural attacks  Wang and Zhang (2023); Lin et al. (2020, 2024); Song et al. (2020); Zhang et al. (2022), which use deep learning to generate realistic profiles.
@@ 5 Conclusion
The main constraints can be summarized in the following two aspects: Firstly, although the black-box text attack model does not require access to the victim model’s parameters and gradients, it necessitates multiple queries to the model. And it is challenging to query the model in real-world large-scale recommendation systems. Secondly, this study solely focuses on the content features of text modality. In reality, recommendation systems encompass other modalities such as images and videos. The issue of attacking models based on these modalities also represents a worthy direction for research.
@@ A.1 Victim Models
  • CoLLM Zhang et al. (2023b). CoLLM seamlessly incorporates collaborative information into LLMs for recommendation by mapping ID embedding to the input token embedding space of LLM.
@@ A.2 Compared Shilling Attacks
  • Heuristic attacks. Heuristic attacks involve selecting items to create fake profiles based on heuristic rules. Random attack Kaur and Goel (2016a) selects filler items randomly while Bandwagon attack Burke et al. (2005a) selects the popular items as users fake preferences, which is white-box as it requires knowledge of the popularity of items, i.e., the training data.
@@ Appendix C Detailed Experimental Results
  • The performance comparison with traditional shilling attacks of TALLRec and CoLLM are shown in 13 and 14.
@@ Appendix D Case Studies
In this section, we present examples of attacks and defenses against RecFormer as a victim model, as shown in Table 15 - 20.
Model Text Exposure
@@
Model Text Exposure
@@
Model Text Exposure
@@
Model Text Exposure
@@
Model Text Exposure
@@
Model Text Exposure

### v2

@@ Abstract
Recently, the powerful large language models (LLMs) have been instrumental in propelling the progress of recommender systems (RS). However, while these systems have flourished, their susceptibility to security threats has been largely overlooked. In this work, we reveal that the introduction of LLMs into recommendation models presents new security vulnerabilities due to their emphasis on the textual content of items. We demonstrate that attackers can significantly boost an item’s exposure by merely altering its textual content during the testing phase, without requiring direct interference with the model’s training process. Additionally, the attack is notably stealthy, as it does not affect the overall recommendation performance and the modifications to the text are subtle, making it difficult for users and platforms to detect. Our comprehensive experiments across four mainstream LLM-based recommendation models demonstrate the superior efficacy and stealthiness of our approach. Our work unveils a significant security gap in LLM-based recommendation systems and paves the way for future research on protecting these systems.  [^1] The source code will be released at https://github.com/CRIPAC-DIG/RecTextAttack.
@@
Jinghao Zhang1,2, Yuting Liu3, Qiang Liu1,2, Shu Wu1,2thanks: To whom correspondence should be addressed., Guibing Guo3, Liang Wang1,2, 1New Laboratory of Pattern Recognition (NLPR), State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS), Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Northeastern University, China jinghao.zhang@cripac.ia.ac.cn, {yutingliu}@stumail.neu.edu.cn {qiang.liu,shu.wu,wangliang}@nlpr.ia.ac.cn, {guogb}@swc.neu.edu.cn
@@ 1 Introduction
Contrary to traditional recommendation models, which rely heavily on abstract and less interpretable ID-based information, LLM-based recommendation models exploit the semantic understanding and strong transferability of LLMs. This approach places a heightened focus on the textual content of items, such as titles and descriptions Lin et al. (2023a); Chen et al. (2023). For instance, many researchers Hou et al. (2022, 2023a); Yuan et al. (2023); Li et al. (2023a); Yang et al. (2023); Geng et al. (2022); Cui et al. (2022); Bao et al. (2023a); Zhang et al. (2023b); Li et al. (2023b); Zhang et al. (2023c) have explored modeling user preferences and item characteristics through a linguistic lens. This methodology promises a revolutionary shift in the conventional paradigm of recommendations by providing generalization capabilities to novel items and datasets.
@@
Despite these advancements, the security of RS remains a largely unaddressed issue. Malicious attacks on these systems can lead to undesirable outcomes, such as the unwarranted promotion of low-quality products in e-commerce platforms or the spread of misinformation in news dissemination contexts. Traditional shilling attack strategies on RS Wang et al. (2023a, 2024c) involve the generation of fake users who are programmed to give high ratings to specific target items. By introducing such cheating data, it aims at influencing the training of the recommender models and subsequently increasing the exposure of the target items.
@@
We construct comprehensive experiments on four mainstream LLM-based recommendation models Geng et al. (2022); Bao et al. (2023a); Li et al. (2023a); Zhang et al. (2023c) as victim models to validate the outstanding efficacy and stealthiness of the textual attack paradigm compared with traditional shilling attacks Burke et al. (2005b); Kaur and Goel (2016); Lin et al. (2020). We further delve into the effects of model fine-tuning and item popularity on the attack. Additionally, we investigate the transferability of the attack across various victim models and recommendation tasks to demonstrate its practical applicability and utility in real-world scenarios. Finally, we evaluate a simple re-writing defense strategy, which also can mitigate the issue to some extent.
@@ 3.1.1 Victim Models
We choose four mainstream LLM-based recommendation models as our victim models: Recformer Li et al. (2023a), P5 Geng et al. (2022), TALLRec Bao et al. (2023a) and CoLLM Zhang et al. (2023c). Please refer to Appendix A.1 for more details.
@@ 3.1.2 Compared Shilling Attacks
Shilling attacks aim to generate fake users that assign high ratings for a target item, while also rating other items to act like normal users for evading. We compare our text attack paradigm with white-box shilling attacks Random attack Kaur and Goel (2016), Bandwagon attack Burke et al. (2005b), and gray-box Aush Lin et al. (2020) and Leg-UP Lin et al. (2024). Please refer to Appendix A.2 for more details.
@@ 3.1.3 Datasets
We conduct experiments on three categories of widely-used Li et al. (2023a); Geng et al. (2022); Bao et al. (2023a); Zhang et al. (2023c) Amazon review dataset introduced by McAuley et al. (2015): ‘Beauty’, ‘Toys and Games’, ’Sports and Outdoors’, which are named as Beauty, Toys and Sports in brief. We use the 5-core version of Amazon datasets where each user and item have 5 interactions at least. The statistics of these datasets are summarized in Appendix A.3.
@@ 3.4 The influence of item popularity
In this section, we examined the impact of the initial popularity of target items on attacks. Popularity bias in recommendation systems favors popular items over personalized ones, limiting diversity, fairness and potentially dissatisfying users’ preferences Zhang et al. (2021b). We selected the top 150 and bottom 150 items in popularity from the dataset as target items and presented the attack results in Table 4. We can observe that items with high popularity experience greater promotion, thereby exacerbating popularity bias. High popularity target items can achieve greater exposure boosts with fewer queries and higher semantic consistency.
@@ 4.1 LLM-based Recommendation
The techniques used by LLMs in the recommendation domain Li et al. (2024); Bao et al. (2023b); Lin et al. (2023b) involve translating recommendation tasks into natural language tasks and adapting LLMs to generate recommendation results directly. These generative LLM-based approaches can be further divided into two paradigms based on whether parameters are tuned: non-tuning and tuning paradigms. The non-tuning paradigm assumes LLMs already have the recommendation abilities and attempt to trigger the strong zero/few-shot abilities by introducing specific prompts Liu et al. (2023); Dai et al. (2023); Mysore et al. (2023); Wang et al. (2023b); Hou et al. (2023b); Wang et al. (2024a). The tuning paradigm uses fine-tuning, prompt learning, or instruction tuning Kang et al. (2023); Bao et al. (2023a); Wang et al. (2022); Geng et al. (2022); Cui et al. (2022) to enhance LLM’s recommendation abilities by using LLMs as encoders to extract user and item representations and then fine-tuning their parameters on specific loss functions.
@@ 4.2 Shilling Attack
Shilling attacks aim to interfere with the recommendation strategy of a victim recommender system by injecting fake users into the training matrix Deldjoo et al. (2019); Toyer et al. (2023); Wang et al. (2024b). This can be implemented through (1) heuristic attacks Burke et al. (2005a); Linden et al. (2003); Kaur and Goel (2016), where fake profiles are created based on subjective inference and existing knowledge; (2) gradient attacks Fang et al. (2020); Li et al. (2016); Zhang et al. (2020); Fang et al. (2018); Huang et al. (2021), which optimize the objective function through a continuous space; and neural attacks  Wang and Zhang (2023); Lin et al. (2020, 2024); Song et al. (2020); Zhang et al. (2022b), which use deep learning to generate realistic profiles.
@@ 5 Conclusion
The main constraints can be summarized in the following two aspects: Firstly, although the black-box text attack model does not require access to the victim model’s parameters and gradients, it necessitates multiple queries to the model. And it is challenging to query the model in real-world large-scale recommendation systems Ying et al. (2018). Secondly, this study solely focuses on the content features of text modality. In reality, recommendation systems encompass other modalities such as images and videos Wei et al. (2019); Zhang et al. (2021a, 2022a, 2023a). The issue of attacking models based on these modalities also represents a worthy direction for research.
@@
Acknowledgements
This work is supported by National Natural Science Foundation of China (62141608, 62206291, 62236010).
@@ A.1 Victim Models
  • CoLLM Zhang et al. (2023c). CoLLM seamlessly incorporates collaborative information into LLMs for recommendation by mapping ID embedding to the input token embedding space of LLM.
@@ A.2 Compared Shilling Attacks
  • Heuristic attacks. Heuristic attacks involve selecting items to create fake profiles based on heuristic rules. Random attack Kaur and Goel (2016) selects filler items randomly while Bandwagon attack Burke et al. (2005b) selects the popular items as users fake preferences, which is white-box as it requires knowledge of the popularity of items, i.e., the training data.
@@ Appendix C Detailed Experimental Results
  • The performance comparison with traditional shilling attacks of TALLRec and CoLLM are shown in 15 and 16.
  • We supplemented the experimental results on the MIND-small news recommendation dataset Wu et al. (2020) with RecFormer as victim model. The performance is shown in Table 9. It can be observed that in the news scenario, LLM-based recommendation systems are also vulnerable to text attacks. This could potentially boost the dissemination of low-quality, fake news, or news with evident bias.
  • We conducted experiments on the impact of different numbers of fake users on the effectiveness of shilling attacks in LLM-based recommendation systems. The results are shown in Table 10. We can observe that as the number of fake users increases, the exposure of the target items also improves. However, compared to the effects of text attacks, this improvement is limited. Additionally, shilling attack shows less stealthiness. As the number of fake users increases, there is a noticeable decline in the overall performance of the recommendation model, which could make it easier for platforms to detect such attacks.
Method Exposure Rel. Impro. # queries
Clean 0.00723 - -
DeepwordBug 0.02655 267.2% 39.6
TextFooler 0.04561 530.8% 101.7
PuncAttack 0.02712 275.1% 60.1
BertAttack 0.04178 477.9% 104.3
Table 9: Performance comparison of attacking finetuned Recformer on MIND-small dataset.
Number of fake users Exposure NDCG@10
0 (Clean) 0.00422 0.03672
100 0.00434 0.03508
200 0.00448 0.03514
500 0.00463 0.03466
1000 0.00481 0.03403
TextFooler 0.01886 0.03596
Table 10: Comparison of different numbers of fake users of shilling attack method Bandwagon on Amazon-Toys dataset (RecFormer as victim model).
@@ Appendix D Case Studies
In this section, we present examples of attacks and defenses against RecFormer as a victim model, as shown in Table 17 - 22.
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure
@@
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure
@@
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure
@@
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure
@@
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure
@@
{tblr}
cells = c, hline1,13 = -0.08em, hline2, 3, 5, 7, 9, 11 = -, Model & Text Exposure